a cylinder 12" by 2" as radius had three tennis balls. how do you calculate the volume of space between the tennis balls?
Answer by mananth(16946)

Volume of cylinder = pir^2h


=pi*2*2*12
48pi
Volume of sphere = 4/3 pi r^3
= 4/3 * pi*2^3
=32/3 *pi
3 Balls
so volume of three balls = 32/3 pi * 3
= 32pi

Volume ofcylinder - volume of 3 balls = vacant space
48pi-32pi
=16pi cubic inches
Ananth